112 students are asked about their plans for the weekend. Of these, 62 plan to attend the home football game, 38 plan to attend the Broadway play at the Auditorium, and 15 plan to attend both. How many plan to attend neither event?
Answer by ikleyn(52835)

.
How many students plan to attend at least one event?

Obviously,  it is  62 + 38 = 100  minus those  15  students who plan to attend both events and whom we counted twice in this sum.  It is  100 - 75 = 85  students.

The rest of  85  students plan to attend neither event.  Their number is  112 - 85 = 27.

Answer.  27 students plan to attend neither event.